OpenStudy (anonymous):

1+sinx/tanx=cosx+cotx help!!!!!!!!!!!

OpenStudy (cggurumanjunath):

check the q ! again @peterchenbmx

OpenStudy (anonymous):

i dont know how they equal out

OpenStudy (cggurumanjunath):

they are not equal !

OpenStudy (cggurumanjunath):

@peterchenbmx

OpenStudy (anonymous):

how do i prove that

OpenStudy (cggurumanjunath):

do u have the q with u ?

OpenStudy (cggurumanjunath):

@peterchenbmx

OpenStudy (anonymous):

to prove: (1+sinx)/tanx=cosx+cotx RHS: cosx + cotx = cosx + 1/tanx = cos x + cosx/sinx = cosx(1+1/sinx) = (cosx/sinx)(sinx + 1) = (1/tanx)(sinx + 1) = (1+sinx)/tanx = LHS